VIBRONIC SPECTROSCOPY OF PHENYLVINYLNITRILE

\underline{DEEPALI N. MEHTA}, POLINA NAVOTNAYA, ALEX PAROBEK, RACHEL CLAY-TON, VANESA VAQUERO VARA, and TIMOTHY S. ZWIER, Department of Chemistry, Purdue University, West Lafayette, IN 47907-2084 U.S.A.

This talk will present results of a gas phase, jet-cooled vibronic spectroscopy study of phenylvinylnitrile (C_{6}H_{5}-CH=CH-CN, PVN). This molecule is seen as a potential photochemical precursor to nitrogen heteroaromatics, and therefore is particularly relevant to Titan’s atmosphere, where nitriles exist in significant abundance. As a first step towards such photochemical studies, a fluorescence excitation spectrum of PVN spanning the range 33,500-35,840 cm^{−1} (298.5-279.0 nm) has been recorded, and dispersed fluorescence spectra that uncover and characterize the vibronic activity have been acquired. The S_{0}-S_{1} origin is a prominent band located at 33,826 cm^{−1}, and the excitation spectrum is characterized by significant vibronic activity leading to spectral congestion. Hole-burning in the region of 33,500-35,840 cm^{−1} has also been completed, and the excitation spectrum can be assigned solely to (E)-PVN. As PVN is the nitrogen-containing counterpart to phenylvinylacetylene (PVA), a comparison of the vibronic spectroscopy between the two molecules will be made.